# Neural Incremental Input-to-State Stable Control Lyapunov Functions for Unknown Continuous-time Systems

### v1

@@ 4.1 Formulation of Loss Functions
The formulation of suitable loss functions is important to train the $\delta$-ISS-CLF and the controller such that minimization of the loss function leads to the satisfaction of the conditions in SCP (9) for the training data set (8) and to the satisfaction of the validity condition (10).
Now, we consider the conditions of the SCP as sub-loss functions to formulate the actual loss function. The sub-loss functions are:
$L_{0}(\psi,\eta)$ $=\sum_{x,\hat{x}\in\mathcal{X}}\max\big{(}0,(-V_{\theta,b}(x,\hat{x})+k_{1}|x-\hat{x}|^{\gamma_{1}}-\eta)\big{)},$ (11a)
$L_{1}(\psi,\eta)$ $=\sum_{x,\hat{x}\in\mathcal{X}}\max\big{(}0,(V_{\theta,b}(x,\hat{x})-k_{2}|x-\hat{x}|^{\gamma_{2}}-\eta)\big{)},$ (11b)
$L_{2}(\psi,\eta)$ $=\sum_{x,\hat{x}\in\mathcal{X},w,\hat{w}\in{\mathcal{W}}}\max\big{(}0,(\frac{\partial V_{\theta,b}}{\partial x}f(x,g_{\bar{\theta},\bar{b}}(x,w))+\frac{\partial V_{\theta,b}}{\partial\hat{x}}f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w}))+\kappa V_{\theta,b}(x,\hat{x})-k_{w}|w-\hat{w}|^{\gamma_{w}}-\eta)\big{)},$ (11c)
$L_{3}(\psi,\eta)$ $=\sum_{x\in\mathcal{X},w\in{\mathcal{W}}}\max\big{(}0,(-\frac{\partial h}{\partial x}\Bigr{\rvert}_{x=x}(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-\mu_{h}h(x)-\eta)\big{)},$ (11d)
where $\psi=[\theta,b,\bar{\theta},\bar{b}]$ and $\eta$ are trainable parameters. As mentioned, the actual loss function is a weighted sum of the sub-loss functions and is denoted by
$L(\psi,\eta)=c_{0}L_{0}(\psi,\eta)+c_{1}L_{1}(\psi,\eta)+c_{2}L_{2}(\psi,\eta)+c_{3}L_{3}(\psi,\eta),$ (12)
where $c_{0},c_{1},c_{2},c_{3}\in{\mathbb{R}}^{+}$ are the weights of the sub-loss functions $L_{0}(\psi,\eta),L_{1}(\psi,\eta),L_{2}(\psi,\eta),L_{3}(\psi,\eta)$, respectively.
@@
$\begin{bmatrix}A\\ B\end{bmatrix}^{\top}\begin{bmatrix}2\alpha\beta\Lambda&-(\alpha+\beta)\Lambda\\ -(\alpha+\beta)\Lambda&2\Lambda\end{bmatrix}\begin{bmatrix}A\\ B\end{bmatrix}+\begin{bmatrix}\mathcal{L}_{L}^{2}\textbf{I}&0&0&0\\ 0&0&0&0\\ 0&0&0&-\theta_{\mathcal{N}}^{\top}\\ 0&0&-\theta_{\mathcal{N}}&\textbf{I}\end{bmatrix}\geq 0,$ (13)
@@
where $\theta_{0},\ldots,\theta_{\mathcal{N}}$ are the weights of the neural network, $\Lambda\in\mathcal{D}_{\geq 0}^{n_{i}},i\in\{1,\ldots,\mathcal{N}\}$, where $n_{i}$ denotes number of neurons in $i$-th layer, and $\alpha,$ and $\beta$ are the minimum and maximum slope of the activation functions, respectively.
@@
The proof can be found in Appendix C. $\square$
@@
The proof shows that the derivative of a neural network with ${\mathcal{N}}$ layers and Lipschitz continuous activation functions is another neural network, where the weights of the derivative network are the same as the original network for ${\mathcal{N}}-1$ layers, and the change is mainly reflected in the weights of the last layer as given in (14b).
Now to ensure the loss function satisfies the matrix inequalities corresponding to the Lipschitzness of Lyapunov, its derivative and the controller, we characterize another loss function denoted by,
$L_{M}(\psi,\Lambda,\hat{\Lambda},\bar{\Lambda})$ $=-c_{l_{1}}\log\det(M_{\mathcal{L}_{L}}(\theta,\Lambda))-c_{l_{2}}\log\det(M_{\mathcal{L}_{dL}}(\hat{\theta},\hat{\Lambda}))-c_{l_{3}}\log\det(M_{\mathcal{L}_{C}}(\bar{\theta},\bar{\Lambda})),$ (15)
where $c_{l_{1}},c_{l_{2}},c_{l_{3}}\in{\mathbb{R}}^{+}$ are weights for sub-loss functions, $M_{\mathcal{L}_{L}}(\theta,\Lambda),M_{\mathcal{L}_{dL}}(\hat{\theta},\hat{\Lambda})$, and $M_{\mathcal{L}_{C}}(\bar{\theta},\bar{\Lambda})$ are the matrices corresponding to the bounds $\mathcal{L}_{L},\mathcal{L}_{dL}$, and $\mathcal{L}_{C}$, respectively.
Finally, to ensure the satisfaction of the validity condition in (10), we propose another loss function:
$L_{v}(\eta)=\mathcal{L}\varepsilon+\eta,$ (16)
where $\mathcal{L}$ is computed as in Theorem 3.2. This completes the characterization of corresponding loss functions for the training algorithm.
@@ 4.2 Training with Formal Guarantee
The training process of the neural Lyapunov function and the controller is described below:
  1. 1.
    Require: Set of sampled data points $\mathcal{X}$ and ${\mathcal{W}}$, and the black box model of the system representing $f(x,w)$.
  2. 2.
    Fix all hyper parameters $\mathcal{L}_{L},\mathcal{L}_{dL},\mathcal{L}_{C},\varepsilon,c=[c_{0},c_{1},c_{2},c_{3},c_{l1},c_{l2},c_{l3}],k=[k_{1},k_{2},k_{w}],\gamma=[\gamma_{1},\gamma_{2},\gamma_{w}]$ a priori. Fix the number of epochs $n_{ep}$ a priori as well.
  3. 3.
    Estimate the Lipschitz constant $\mathcal{L}_{x}$ and $\mathcal{L}_{u}$ using the reverse Weibull distribution [37]. Get the Lipschitz constant $\mathcal{L}$ as defined in Theorem 3.2.
  4. 4.
    Randomly generate several batches from the dataset. Set the number of batches a priori.
  5. 5.
    In each epoch, pass each batch to the networks and get the outputs $\frac{\partial V_{\theta,b}}{\partial x},f(x,g_{\bar{\theta},\bar{b}}(x,w)),\frac{\partial V_{\theta,b}}{\partial\hat{x}},f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w}))$ and $V_{\theta,b}(x,\hat{x})$. Calculate the loss for each batch using the loss functions mentioned in (12), (15) and (16). The cumulative sum of batch losses will give the epoch loss.
  6. 6.
    Utilizing Adam or Stochastic Gradient Descent (SGD) optimization techniques with a specified learning rate, reduce the loss function and update the trainable parameters $\phi,\eta$. The learning rate can be different for different parameters.
  7. 7.
    Repeat steps 5 and 6 until the loss functions converge according to Theorem 4.4 defined next. After successful convergence of the training, the neural network will act as $\delta$-ISS control Lyapunov function $V_{\theta,b}$ under the action of the controller $g_{\bar{\theta},\bar{b}}$, and the closed-loop system is assured to be incrementally input-to-state stable.
Finally, we present the theorem that provides the formal guarantee for the incremental stable nature of the closed-loop system under the action of the controller $g_{\bar{\theta},\bar{b}}$.
@@
Consider a continuous-time control system $\Xi$ as in (1) with compact state-space ${\mathbf{X}}$ and input space ${\mathbf{W}}$. Let, $V_{\theta,b}$ and $g_{\bar{\theta},\bar{b}}$ denote the trained neural networks representing the $\delta$-ISS control Lyapunov function and the controller such that $L(\psi,\eta)=0,L_{v}(\eta)=0$ and $L_{M}(\psi,\Lambda,\hat{\Lambda},\bar{\Lambda})\leq 0$ over the training data sets $\mathcal{X}$ and $\mathcal{W}$. Then, the closed-loop system under the influence of the controller $g_{\bar{\theta},\bar{b}}$ is guaranteed to be incrementally ISS within the state-space.
@@
The proof can be staged over satisfying conditions for the loss functions.
  • The first loss $L(\psi,\eta)=0$ implies that the SCP has been solved with the minimum value being $\eta_{S}^{*}$. Hence, the controller is trained to ensure that the closed-loop system is $\delta$-ISS with initial conditions belong to $\mathcal{X}$ (that is, finite over data).
  • The second loss $L_{v}(\eta)=0$ implies the satisfaction of Theorem 3.2, which means that the solution of SCP is also valid for the ROP (6). Hence, the closed-loop system is $\delta$-ISS for any initial state in ${\mathbf{X}}$ and any external input sequence ${\mathsf{w}}$.
  • The third loss $L_{\mathcal{M}}(\psi,\Lambda,\hat{\Lambda},\bar{\Lambda})\leq 0$ implies that the matrices $M_{\mathcal{L}_{L}}(\theta,\Lambda),M_{\mathcal{L}_{dL}}(\hat{\theta},\hat{\Lambda})$ and $M_{\mathcal{L}_{C}}(\bar{\theta},\bar{\Lambda})$ are positive definite, satisfying the Lipschitz continuity assumption of the neural networks with the assumed Lipschitz bounds.
Hence, the satisfaction of the above theorem leads to ensuring that the closed-loop system is $\delta$-ISS under the action of the controller. This completes the proof. $\square$
@@
If the algorithm does not converge successfully, one cannot judge the incremental input-to-state stability of the system with the specified hyperparameters $c,k,\gamma,\mathcal{L}_{L},\mathcal{L}_{dL},\mathcal{L}_{C}$ and reconsider the initial parameters.
@@
In addition, the initial feasibility of condition (13) is required to satisfy the criterion of loss $L_{M}$ in (15) according to Theorem 4.4. Choosing small initial weights and bias of the neurons can ensure this condition [26].
Remark 4.7
The class $\mathcal{KL}$ functions of Definition 2.5 can be computed using a similar procedure once the trained $\delta$-ISS-CLF is obtained as referred in [40]. The methodology of obtaining the functions will follow the proof mentioned in A.
@@ 5.1 Scalar Nonaffine Nonlinear System
where $\mathsf{x}(t)$ denotes the state of the system at time instant $t$. The constant $a$ represents the rate constant of the system. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{2},\frac{\pi}{2}]$. Moreover, the external input set is bounded within ${\mathbf{W}}=[-0.5,0.5]$. We consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=0.2,\mathcal{L}_{u}=0.25$ using the results in [37].
@@
As can be seen in Figure 1(a), under the different input conditions, the trajectories starting from different initial conditions maintain the same distance after some time instances under the influence of the controller. The $\delta$-ISS-CLF plot for this case is shown in Figure 1(b). Note that in Figure 1(c), the plot of condition (6c) is shown under the same input signal, and it is always negative, satisfying the condition.
Figure 1: (a)Trajectories starting from different initial conditions under different input signals (for all $t\geq 0,{\mathsf{w}}(t)=-0.1\in{\mathbf{W}}$ for blue curve and ${\mathsf{w}}(t)=0.1\in{\mathbf{W}}$ for red curve), (b) The $\delta$-ISS-CLF plot is greater than zero for all $(x,\hat{x})\in{\mathbf{X}}\times{\mathbf{X}}$, (c) The derivative of $V$ is always negative corresponding to the same input ${\mathsf{w}}=\hat{{\mathsf{w}}}$ satisfying the condition (6c).
@@ 5.2 Single-link Manipulator
We consider a single link manipulator dynamics [21, 24], whose dynamics is governed by the following set of equations.
@@
where $\mathsf{x}_{1}(t),\mathsf{x}_{2}(t)$ denotes angular position and velocity at time instance $t$ respectively. The constants $M=1,b=0.1$ represent the mass and damping coefficient of the system, respectively, while $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{6},\frac{\pi}{6}]\times[-\frac{\pi}{6},\frac{\pi}{6}]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.5,0.5]$. Also, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=1.005,\mathcal{L}_{u}=0.01$.

### v2

@@ 4.1 Formulation of Loss Functions
The formulation of suitable loss functions is important to train the $\delta$-ISS-CLF and the controller such that minimization of the loss function leads to the satisfaction of the conditions (9) for the training data set (7). Now, we consider the conditions as sub-loss functions mentioned below:
$L_{1}(\psi)$ $),\displaystyle=\sum_{x,\hat{x}\in\mathcal{X}}\max\big(0,(V_{\theta,b}(x,\hat{x}$$\big)$, (10a)
$L_{2}(\psi)$ $),\displaystyle=\sum_{x,\hat{x}\in\mathcal{X}}\max\big(0,(-V_{\theta,b}(x,\hat{x})+k_{1}|x-\hat{x}|^{\gamma_{1}}-\hat{\eta}$$\big)$, (10b)
$L_{3}(\psi)$ $),\displaystyle=\sum_{x,\hat{x}\in\mathcal{X}}\max\big(0,(V_{\theta,b}(x,\hat{x})-k_{2}|x-\hat{x}|^{\gamma_{2}}-\hat{\eta}$$\big)$, (10c)
$L_{4}(\psi)$ $={\hskip-14.22636pt\sum_{x,\hat{x}\in\mathcal{X},w,\hat{w}\in{\mathcal{W}}}\hskip-14.22636pt\max\big(0,(\widehat{\mathsf{L}}_{f,g}V(x_{q},x_{r})+\kappa V_{\theta,b}(x,\hat{x})-k_{w}|w-\hat{w}|^{\gamma_{w}}+\delta_{V}-\hat{\eta})\big),}$ (10d)
$L_{5}(\psi)$ $={\hskip-14.22636pt\sum_{x\in\mathcal{X},w\in{\mathcal{W}}}\hskip-11.38092pt\max\big(0,(-\widehat{\mathsf{L}}_{f,g}h(x_{q})-\mu_{h}h(x_{q})-\delta_{h}-\hat{\eta})\big)}$, (10e)
where $\psi=[\theta,b,\bar{\theta},\bar{b}]$ are trainable parameters. As mentioned, the actual loss function is a weighted sum of the sub-loss functions and is denoted by
$L(\psi)=\sum_{i=1}^{5}c_{i}L_{i}(\psi),$ (11)
where $c_{i}\in{\mathbb{R}}^{+},i\in[1;5]$ are the weights of the sub-loss functions $L_{i}(\psi)$.
@@
$\begin{bmatrix}A\\ B\end{bmatrix}^{\top}\begin{bmatrix}2\alpha\beta\Lambda&-(\alpha+\beta)\Lambda\\ -(\alpha+\beta)\Lambda&2\Lambda\end{bmatrix}\begin{bmatrix}A\\ B\end{bmatrix}$+
$\begin{bmatrix}\mathcal{L}_{L}^{2}\textbf{I}&0&0&0\\ 0&0&0&0\\ 0&0&0&-\theta_{\mathcal{N}}^{\top}\\ 0&0&-\theta_{\mathcal{N}}&\textbf{I}\end{bmatrix}\geq 0,$ (12)
@@
where $\theta_{0},\ldots,\theta_{\mathcal{N}}$ are the weights of the neural network, $\Lambda\in\mathcal{D}_{\geq 0}^{n_{i}},i\in\{1,\ldots,\mathcal{N}\}$, where $n_{i}$ denotes number of neurons in $i$-th layer, and $\alpha,$ and $\beta$ are the minimum and maximum slope of the activation functions, respectively. It should be carefully noted that the dimensions of the matrices $A$ and $B$ are $(\sum_{i=1}^{\mathcal{N}}n_{i})\times(1+\sum_{i=0}^{\mathcal{N}-1}n_{i})$ and $(\sum_{i=1}^{\mathcal{N}}n_{i})\times(1+\sum_{i=1}^{\mathcal{N}}n_{i})$ respectively
@@
The proof can be found in Appendix F. $\square$
@@
The proof shows that the derivative of a neural network with $\mathcal{N}$} layers and Lipschitz continuous activation functions is another neural network, where the change is mainly reflected in the weights of the last layer as given in (13b).
Now, to ensure the loss function satisfies the matrix inequalities corresponding to the Lipschitzness of Lyapunov, its derivative and the controller, we characterize another loss function denoted by,
$L_{M}(\psi,\Lambda,\hat{\Lambda},\bar{\Lambda})=-c_{l_{1}}\log\det(M_{\mathcal{L}_{L}}(\theta,\Lambda))$
$-c_{l_{2}}\log\det(M_{\mathcal{L}_{dL}}(\hat{\theta},\hat{\Lambda}))-c_{l_{3}}\log\det(M_{\mathcal{L}_{C}}(\bar{\theta},\bar{\Lambda})),$ (14)
[…]
@@ 4.2 Training with Formal Guarantee
The training process of the neural Lyapunov function and the controller is described in Algorithm 1.
Algorithm 1 Training of the Neural Networks
1:Black box model of the system $f(x,u)$, Data sets: $\mathcal{X},\mathcal{W}$
[…]
3:Select the hyperparameters $\epsilon$, $c=[c_{0},c_{1},c_{2},c_{3},c_{l1},c_{l2},c_{l3}],k=[k_{1},k_{2},k_{w}],\Upsilon=[\gamma_{1},\gamma_{2},\gamma_{w}],\kappa,\mu_{h},\mathcal{L}_{L},\mathcal{L}_{C},\mathcal{L}_{dL},\mathcal{L}_{h}$ and number of epochs $n_{ep}$.
4:Computation of Lipschitz constants $\mathsf{L}_{1},\mathsf{L}_{2}$ and $\mathsf{L}_{u}$. Compute $\mathcal{L}$ using Theorem 3.5.
5:Initialize Neural networks and trainable parameters $\theta,b,\bar{\theta},\bar{b},\Lambda,\bar{\Lambda}$. Initialize $\hat{\eta}$ as $\hat{\eta}=-\mathcal{L}\varepsilon$.
6:for $i\leq Epochs$ (Training starts here) do
7:  Create batches of training data from $\mathcal{X},\mathcal{W}$
8:  Find batch losses using (11) and (4.1).
9:  Use ADAM or SGD optimizer with specified learning rate [ruder2016] to reduce loss and update the trainable parameters.
10:  If Theorem 4.4 is satisfied $\rightarrow$ break.
11:end for
12:return $V_{\theta,b},g_{\bar{\theta},\bar{b}}.$
We present the theorem that provides the formal guarantee for the incremental stable nature of the closed-loop system under the action of the controller $g_{\bar{\theta},\bar{b}}$.
@@
Consider a continuous-time control system $\Xi$ as in (1) with compact state-space $\mathbf{X}$} and input space $\mathbf{W}$}. Let, $V_{\theta,b}$ and $g_{\bar{\theta},\bar{b}}$ denote the trained neural networks representing the $\delta$-ISS control Lyapunov function and the controller such that $L(\psi,\eta)=0$ and $L_{M}(\psi,\Lambda,\hat{\Lambda},\bar{\Lambda})\leq 0$ over the training data sets $\mathcal{X}$ and $\mathcal{W}$. Then, the closed-loop system under the influence of the controller $g_{\bar{\theta},\bar{b}}$ is guaranteed to be incrementally ISS within the state-space.
@@
The first loss $L(\psi,\eta)=0$ is defined as a weighted sum of its constituent sub-loss terms, therefore the condition $L(\psi,\eta)=0$ implies that its individual sub-loss components vanishes, which in turn, ensures the conditions of Theorem 3.5 has been solved with $\hat{\eta}$. This enforces the satisfaction of all the inequalities of the $\delta$-ISS-CLF conditions corresponding to Definition 2.5 for the sampled datapoints in $\mathcal{X}$ and $\mathcal{W}$ corresponding to (7). Note that with the initialization of $\hat{\eta}=-\mathcal{L}\varepsilon$, the condition of Theorem 3.5 has already been satisfied, resulting closed-loop system to be $\delta$-ISS for any initial state in $\mathbf{X}$} and any external input $\mathbf{W}$}. Another loss $L_{\mathcal{M}}(\psi,\Lambda,\hat{\Lambda},\bar{\Lambda})\leq 0$ ensures the predefined Lipschitz constants of the networks. Hence, the satisfaction of the above theorem leads to ensuring that the closed-loop system is $\delta$-ISS under the action of the controller. $\square$
@@
Note that if the algorithm fails to converge, the $\delta$-ISS of the closed-loop system cannot be determined for the given hyperparameters $c,k,\Gamma,\mathcal{L}_{L},\mathcal{L}_{C}$. Convergence can be improved by reducing the discretization parameter $\epsilon$ [zhao2021learning], tuning neural network hyperparameters (e.g., architecture, learning rate) [nn_lr], or revising the algorithm’s initial hyperparameters.
@@
In addition, the initial feasibility of condition (4.1) is required to satisfy the criterion of loss $L_{M}$ in (4.1) according to Theorem $4.$ Choosing small initial weights and bias of the neurons can ensure this condition [pauli2022a].
@@ 5.1 Scalar Nonaffine Nonlinear System
where $\mathsf{x}(t)$ denotes the state of the system at time instant $t$. The constant $a$ represents the rate constant of the system. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{2},\frac{\pi}{2}]$. Moreover, the external input set is bounded within ${\mathbf{W}}=[-0.5,0.5]$. We consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=0.2,\mathcal{L}_{u}=0.25$ using the results in [Lipschitz].
@@
As can be seen in Figure 1(a), under the different input conditions, the trajectories starting from different initial conditions maintain the same distance after some time instances under the influence of the controller. The $\delta$-ISS-CLF plot for this case is shown in Figure 1(b). Note that in Figure 1(c), the plot of condition (6d) is shown under the same input signal, and it is always negative, satisfying the condition.
Figure 1: (a)Trajectories starting from different initial conditions under different input signals (for all $t\geq 0,{\mathsf{w}}(t)=-0.1\in{\mathbf{W}}$ for blue curve and ${\mathsf{w}}(t)=0.1\in{\mathbf{W}}$ for red curve), (b) The $\delta$-ISS-CLF plot is greater than zero for all $(x,\hat{x})\in{\mathbf{X}}\times{\mathbf{X}}$, (c) The derivative of $V$ is always negative corresponding to the same input $^{\mathsf{w}}=\hat{{\mathsf{w}}}$ satisfying the condition (6d).
@@ 5.2 Single-link Manipulator
We consider a single link manipulator dynamics [lewis2020neural, murray2017mathematical], whose dynamics is governed by the following set of equations.
@@
where $\mathsf{x}_{1}(t),\mathsf{x}_{2}(t)$ denotes angular position and velocity at time instance $t$ respectively. The constants $M=1,b=0.1$ represent the mass and damping coefficient of the system, respectively. We consider the state space of the system to be ${\mathbf{X}}=[-\frac{\pi}{6},\frac{\pi}{6}]\times[-\frac{\pi}{6},\frac{\pi}{6}]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-0.5,0.5]$. Also, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=1.005,\mathcal{L}_{u}=1$.
@@
5.3 Magnetic Levitator System
For the next case study, we consider a magnetic levitator system [ghosh2014design], whose dynamics is given by:
$\dot{\mathsf{x}}_{1}(t)$ $=\mathsf{x}_{2}(t),$
$\dot{\mathsf{x}}_{2}(t)$ $=g-\frac{k}{m}\frac{i^{2}}{\mathsf{x}_{1}^{2}(t)}$ (17)
where $\mathsf{x}(t)=[\mathsf{x}_{1}(t),\mathsf{x}_{2}(t)]$ denotes the position and velocity of the ball at time instant $t$, $m=0.01$ stands for the mass of the ball, $g=9.81$ being the gravitational constant and $k=0.009$ being the coil constant of the system. The input to the system is the coil current $i$. We consider the state space to be ${\mathbf{X}}=[0.25,0.75]\times[-0.25,0.25]$. For this case study, we show that the controller is able to make the system $\delta$-GAS, as there is no external input. We estimate the Lipschitz constants to be $\mathcal{L}_{x}=1,\mathcal{L}_{u}=0.15$.
We first fix the training hyper-parameters as $\epsilon=0.005,\mathcal{L}_{L}=1,\mathcal{L}_{dL}=1,\mathcal{L}_{C}=2,k=[0.00001,0.2,0.01],\kappa=0.00001,\mu_{h}=0.1$. So, the Lipschitz constant according to Theorem 3.5 is $2.8685$. We fix the structure of $V_{\theta,b}$ as $l_{f}=2,h_{f}^{1}=60$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=20$. The activation functions for $\delta$-ISS-CLF and the controller are Softplus and ReLU, respectively. The training algorithm converges to obtain $\delta$-ISS-CLF $V_{\theta,b}$ along with $\eta=-0.0162$. Hence, $\eta+\mathcal{L}\epsilon=-0.0162+2.8685\times 0.005=--0.00186$, therefore, using Theorem 4.4, we can guarantee that the obtained $\delta$-ISS-CLF $V_{\theta,b}$ is valid and that the closed-loop system is assured to be incrementally input-to-state stable under the influence of the controller $g_{\bar{\theta},\bar{b}}$.
One can see from Figure 3 that under the different input conditions, the trajectories corresponding to various states starting from different initial conditions converge to a particular trajectory after some time instances under the influence of the controller.
Figure 3: Top: Position (left) and velocity (right) of the ball, where the trajectories converge towards each other without the presence of external input for all $t\geq 0$. Bottom: The difference in positions (left) and velocities (right) subjected to different initial conditions and input currents.